20180130 UT E and D night:

OSA: G. Bechetti
ISp: A. Cavallaro
ISA: B. Rothberg

MODS1+2 transform data collection with pointing mask and MODS2 RDG correct RZP

"Movies" which step through all of the transform data taken with MODS1R and MODS2B are attached. These show the position of the mask on the detector remains fixed, but the position of the on-axis star changes from one image to the next, as the guide probe is moved from one to another off-axis guide star. The movies show 3 sequences, each with a different on-axis star, but all with the same position angle.



UT 07:30 - authorized LUCI-2 monocular - handover from INAF
UT 07:37 - pointing and collimation check - seeing is highly variable in DIMM, 1-1.5\x94+, going to
backup LUCI-2 program RJ04 Mergers first G200 HK, will assess seeing before continuing to
zJ. First target is telluric for UGC 5101
UT 07:41 - sending preset for HD 56385 Telluric
Of course\x85 there\x92s two stars.. . but thanks to SIMBAD got the right one.
UT 07:49 - starting integration - guider 0.70\x94 - you\x92re killing me!
- 0016-0018 - Doing \xBE of the telluric exposures, will load zJ telluric and do \xBE then move to
target
- 0019-0021 zJ G200 Telluric
UT 08:00 - moving to UGC 5101 - guider average 0.65\x94
UT 08:10 - starting science integrations
UT 08:13 -error commiting filter wheel2 - requested HKspec it went J-filter, manually put in
HKspec, same error, - motor control did not complete in expected time, then do skip to
spectroscopy config and went, now OK
UT 08:16 - another instance of script pausing - manually clicked GO script continued
UT 08:23 - getting excellent S/N -> will stop early - 15min shutter time
UT08:32 - 0026-0031 - completed UGC 5101 HK only - making an attempt on Dual AGN now
to take advantage of unexpected good seeing
========================================================================
=
-moving to SDSSJ084905+111450 - 1.5\x94 wide slit
#0039-0050 (24 min total)
UT 08:37 - guider 0.75\x94 avg
UT 08:40 - FM4 moved +110 pixels in Y -had turned off flexure comp before slew - went from
+61 dec to +11 dec
UT 08:43 - will shift mirror back - 1039, 121 to 1037, 1
0035 - will be field centered image and to verify FM4 moved back - use 0033 as background for
pair subtraction
-movement ok, now putting mask back in FPU - update IT 5645 (monitor for FM4 motions)
0036 - thru slit
UT 08:50 - DIMM 0.76\x94 guider avg 0.75\x94
UT 08:53 - the \x931024\x94 auto clicked on after restarting the RTD - so it offset to 1024, need to
manually move down DETY 504 pixels or -60\x94
0038 - confirmation thru-slit
UT 08:58 - confirmed both nuclei are well separated - starting science exposures
-seeing 0.68\x94 on DIMM Guider 0.65\x94!
UT 09:02 - brilliant! I can see the fainter 2nd trace 9pixels below the brigher nucleus!!!!!
-Woohoo!!!!
UT 09:11 - airmass 1.172 - seeing 0.68\x94 avg Guider
UT 09:29 - finished Dual AGN, moving to Telluric HIP 43018
#0054-0057 Telluric spectra
UT 09:41 competed Telluric for Dual AGN - LUCI-2 in safe config
========================================================================
=
MODS1&2 Transform Stone I field -data directory is 20180130/ldg /rdg
UT 09:46 - Switching PCSInstrument.conf to updated LDG and RDG
UT 09:50 switching to Binocular MODS for AGW Transform
UT 10:04 -bouncing PCS
UT 10:19 - successfully sent a bincolular preset from IDL, key was to do it all in one window
(might want to add that to wiki page)
UT 10:37 - little snafu, but going - the issue was an errant space in the catalog name \x93Stone_I \x93
vs \x93Stone_I\x94 - going now
-first list is Stone_I_star2_pa135.list
0002 blue MODS2
0003 red MODS1
0004_BG red mods1, 0003_BG mods2 blue - Background image
DIMM 0.81\x94
UT 10:56 - A. Rakich contacted remote room - running up lasers to measure mirrors - will call if
we use LUCI-2 or when we close up to do LUCI-2 cals
UT 11:07 -starting star6 list both sides
-had to manually send a preset on the right side - i initially violated co-pointing limit when i sent
the gs6 list on both sides
-resent gs6 list on right - works OK
UT 11:34 - stars seem to moving within the box (the mask is stationary)
UT 11:41 - left side finished, waiting for right side to finish, will send binocular preset to next
GS8
UT 11:45 - star8 list - sending binocular preset
UT 11:47 - starting data collection star8 list
UT 12:19 - contacted A. Rakich to turn off laser - will be switching back to LUCI-2 given that
switching to LBC and starting up TMS + setup would take us past astronomical twilight.
UT 12:20 - MODS-1 finished
UT 12:21 - MODS-2 finished
Done with AGW Transform
========================================================================
=
UT 12:22 - switch to LUCI-2, putting MODS-1 and MODS-2 to sleep, returning
PCSInstrument.conf to earlier version
UT 12:35 - configured for LUCI-2 - doing a pointing and collimation check
-moving to target for Olga - QSO
UT 12:44 - moving to target - 0.86\x94DIMM
-in acquisition field guider avg 0.65\x94
UT 12:55 - starting science exposure
UT 13:13 - science target finished, moving to Telluric star
UT 13:20 - field image looks out of focus, moving out of FPU, switching to N1.8 and back to
N375
UT 13:23 - accidentally left in N1.8 and took exposure, but looks out of focus #0074
UT 13:24 - put N3.75 in, image looks better, slightly elongated but not diffuse center
UT 13:25 - 0077-0080 telluric spectra
UT 13:34 - closing to take cals
========================================================================
=
UT 13:44 - starting 1\x94 G200 HK cals
UT 13:46 - script jumped all the way to last obs item (6th in the script) - after loading and clicking
GO - last item would have lamp turned on - but it didn\x92t go on - hit reset, this time script worked
0083-0101
UT 13:56 - starting 0.5\x94 G200 HK Cals
-Neon light didn\x92t go on? IT 6943..
0102-0117
0118-0119 -> taking another set of Neon lamps - confirmed Neon didn\x92t go on - Yay! Can use
those as my darks!
UT 14:11 - starting 1.5\x94 G200 HK Cals
0120-0137 (0130-0131 - Neon but not functioning so use these as darks) - this time it went on!)
UT 14:25 - Cals complete - putting LUCI-2 in safe config

-- %USERSIG{OlgaKuhn - 2018-04-27}%

Comments

 
Topic attachments
I Attachment Action Size Date Who Comment
mods1r_all.mpegmpeg mods1r_all.mpeg manage 936 K 27 Apr 2018 - 21:40 OlgaKuhn series of images for mods2r transform measurement
mods2b_all.mpegmpeg mods2b_all.mpeg manage 1 MB 27 Apr 2018 - 21:39 OlgaKuhn series of images for mods2b transform measurement
Topic revision: r1 - 27 Apr 2018, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback